# Robust Semiparametric Graphical Models with Skew-Elliptical Distributions

### v2

@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
Proof.
We have: $\hat{S}^{\rho,\alpha}_{ij}=$ $4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)$ and we introduce $\Sigma^{\rho}$, the correlation matrix using Spearman’s rho statistic. We know that $A_{\rho}=L^{\top}ML$, where $M=I-D^{2}$. By Definition: $\Sigma^{\rho}=L^{\top}D_{\kappa}L$. This implies:
$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
$\left\lVert\widehat{\Sigma}^{\rho}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\rho}}\right\rVert_{\max},$
$\left\lVert 4\sin\left(\frac{\pi}{6}\widehat{\Sigma}^{\rho}\right)\right\rVert_{\max}\geq\left\lVert 4\sin\left(\frac{\pi}{6}A_{\widehat{\rho}}\right)\right\rVert_{\max}.$
[…]
$E(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right].$ (20)
We rewrite (20) as:
$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right),\quad\text{where }a_{ij}=\frac{\pi E(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$
The term $|a_{ij}|\leq\frac{\pi}{n-2}$. Therefore, for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$, and thus:
$\text{pr}\left(\left\lVert 4\sin\left(\frac{\pi}{6}\hat{\Sigma}^{\rho}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)$ $=\text{pr}\left(\max_{ij}\left|4\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=\text{pr}\left(\max_{ij}\left|2\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
$\leq\text{pr}\left(\max_{ij}\frac{2\pi}{3}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})-a_{ij}\right|>\frac{t}{2}\right)\quad\text{(Lipshitz function)}$
$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-|a_{ij}|\right)$
$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-\frac{t}{6}\right)$
$=\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)$
$\leq\sum_{ij}\text{pr}\left(\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)\quad\text{(Boole's inequality)}$
$\leq p^{2}\exp\left(-\frac{nt^{2}(9-2\pi)^{2}}{(2\pi-3)^{2}}\right)\quad\text{(Hoeffding's inequality)}$
Setting:
$\beta=p^{2}\exp\left(\frac{(9-2\pi)^{2}nt^{2}}{(2\pi-3)^{2}}\right)$
and solving for $t$, we get:
$t=\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left(\frac{p^{2}}{\beta}\right)}$
When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left({p^{3}}\right)}$. ∎
@@
Proof.
Similarly to the case of Spearman’s rho, we have that $\hat{S}^{\tau,\alpha}_{ij}=$ $2\sin\left(\frac{\pi}{2}A_{\hat{\tau}_{ij}}\right)$. We introduce $\Sigma^{\tau}$, the correlation matrix using Kendall’s tau statistic. We know that $A_{\tau}=L^{\top}ML$, where $M=I-D^{2}$. By Definition: $\Sigma^{\tau}=L^{\top}D_{\kappa}L$. This implies that:
$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
$\left\lVert\widehat{\Sigma}^{\tau}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\tau}}\right\rVert_{\max},$
$\left\lVert 2\sin\left(\frac{\pi}{2}\widehat{\Sigma}^{\tau}\right)\right\rVert_{\max}\geq\left\lVert 2\sin\left(\frac{\pi}{2}A_{\widehat{\tau}}\right)\right\rVert_{\max}.$
[…]
$E(\hat{\tau}_{ij})=\tau_{ij},$
where
$\tau_{ij}:=\operatorname{Corr}(\operatorname{sign}(X_{i}-\widetilde{X}_{i}),\operatorname{sign}(X_{j}-\widetilde{X}_{j}))$
is the population version.
$\text{pr}\left(\left\lVert 2\sin\left(\frac{\pi}{2}\hat{\Sigma}^{\tau}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)\quad$ $=\text{pr}\left(\max_{ij}\ \left|2\sin\left(\frac{\pi}{2}\hat{\tau}_{ij}\right)-\sin\left(\frac{\pi}{2}\tau_{ij}\right)\right|>t\right)$
$\leq\text{pr}\left(\max_{ij}\pi\left|\hat{\tau}_{ij}-\tau_{ij}\right|>t\right)\quad\text{(Lipshitz function)}$
$\leq\sum_{ij}\text{pr}\left(\left|\hat{\tau}_{ij}-\tau_{ij}\right|>\frac{t}{\pi}\right)\quad\text{(Boole's inequality)}$
$\leq p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right)\quad\text{(Hoeffding's inequality)}$
Setting
$\beta=p^{2}\exp\left(-\frac{nt^{2}}{2\pi^{2}}\right).$
And solving for $t$, we get:
$t=\sqrt{\frac{2\pi^{2}\log\left(\frac{2p^{2}}{\beta}\right)}{n}}.$
When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{2\pi^{2}}{n}\log\left({p^{3}}\right)}.$
A.5 Proof of Theorem 5.3
We start by proving $\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ and $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$. Kendall’s tau and Spearman’s rho statistics can be rewritten in terms of U-statistics, and their estimates of correlation matrix achieve $O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$ as rate of convergence. Under assumption (a), we know that $\Sigma^{0}$ has bounded eigenvalues, which implies:
$\|\hat{L}-L\|_{\max}\leq\tau_{2}\|\hat{\Sigma}-\Sigma^{0}\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right).$
Regarding the convergence of the estimated $\hat{M}$ matrix, we know that it has diagonal elements defined as: $\hat{m}_{kk}=1-\frac{2}{\pi}\frac{\hat{\alpha}_{kk}}{1+\hat{\alpha}_{kk}}$ where $\hat{\alpha}$ parameters are estimated via the EM algorithm introduced in Abe et al., (2021), for which we assume that the regulation conditions of Wu, (1983) hold. Then $\|\hat{M}-M\|_{\max}O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$. We find bounds for $\|\hat{\Omega}-\Omega\|_{\max}$.
$\|\hat{\Omega}-\Omega\|_{\max}$ $=\|2\sin\left(\frac{\pi}{2}(\hat{L}^{\top}\hat{M}\hat{L}^{\top})\right)^{-1}-2\sin\left(\frac{\pi}{2}({L}^{\top}{M}{L}^{\top})\right)^{-1}\|_{\max}$
$\leq\|(\hat{L}^{\top}\hat{M}\hat{L})^{-1}-({L}^{\top}{M}{L})^{-1}\|_{\max}$
$=\|(\hat{L}\hat{M}^{-1}\hat{L}^{\top})-({L}{M}^{-1}{L}^{\top})\|_{\max}$
$=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}+L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}+LM^{-1}(\hat{L}-L)\|_{\max}$
We define $A_{1}=(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}$, $A_{2}=L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}$, and $A_{3}=LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})$.
$\|A_{1}\|_{\max}=\|(\hat{L}-L)\hat{M}^{-1}\hat{L}^{\top}\|_{\max}\leq\|\hat{L}-L\|_{\max}\|\hat{M}^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
$\|\hat{L}-L\|_{\max}=O_{P}\left(\sqrt{\frac{\log p}{n}}\right),\quad\|\hat{M}^{-1}\|_{\max}=O(1),\quad\|\hat{L}^{\top}\|_{1}=O(s).$
Where $s$ are the non-zero elements in the $L$ matrix.
$A_{2}=\|L(\hat{M}^{-1}-M^{-1})\hat{L}^{\top}\|_{\max}\leq\|L\|_{\infty}\|\hat{M}^{-1}-M^{-1}\|_{\max}\|\hat{L}^{\top}\|_{1}$
Where $\|\hat{M}^{-1}-M^{-1}\|_{\max}=\max_{j}\left|\frac{1}{\hat{M}_{jj}}-\frac{1}{M_{0,jj}}\right|=\left|\frac{\hat{M}_{jj}-M_{0,jj}}{\hat{M}_{jj}M_{0jj}}\right|$. This means that: $\|\hat{M}^{-1}-M^{-1}\||_{\max}=O_{P}(\|\hat{M}-M\|_{\max})=O_{P}\left(\sqrt{\frac{\log p}{n}}\right)$.
$\|A_{3}\|_{\max}$ $=\|LM^{-1}(\hat{L}^{\top}-L^{\top}_{0})\|_{\max}$
$\leq\|L\|_{\infty}\|M^{-1}\|_{\infty}\|\hat{L}^{\top}-L^{\top}_{0}\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right)$
We conclude that:
$\|\hat{\Omega}-\Omega\|_{\max}=O_{P}\left(s\sqrt{\frac{\log p}{n}}\right).$
A.6 Proof of Theorem 5.4
As discussed in Lam and Fan, (2009), we only need to prove a bound in $\|\hat{\Omega}-\Omega\|_{\max}$ for the three shrinkage procedures proposed to recovery $\Omega$. For the glasso algorithm, the proof follows the proof of Theorem 2, in Lam and Fan, (2009). For the CLIME, the bounds are provided in Theorem 4 in Cai et al., (2011), under the Frobenius norm, which implies $\|\cdot\|_{\max}$. And for the graphical Dantzig selector, the bounds are provided in Section 3 of Yuan, (2010) work.

### v3

@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
Let $B_{ij}=\dfrac{1+\alpha_{i}\alpha_{j}}{\sqrt{(1+\alpha_{i}^{2})(1+\alpha_{j}^{2})}}$ and, for $i\neq j$, we write the target as $S^{\rho,\alpha}_{ij}=2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\,B_{ij}$, and $S^{\rho,\alpha}_{ii}=1$. For $i\neq j$:
$\left|\widehat{S}^{\rho,\alpha}_{ij}-{S}^{\rho,\alpha}_{ij}\right|$ $\leq\left|\widehat{B}_{ij}2\sin\!\big(\tfrac{\pi}{6}\widehat{\rho}_{ij}\big)-B_{ij}2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\right|$
$\leq|\widehat{B}_{ij}|\left|2\sin\!\big(\tfrac{\pi}{6}\widehat{\rho}_{ij}\big)-2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\right|+|\widehat{B}_{ij}-B_{ij}|,$
since, by construction, $|\widehat{B}_{ij}|\leq 1$. From liu2012high, we know that, with probability $1-1/p^{2}$, we have:
$\max_{ij}\left|2\sin\!\big(\tfrac{\pi}{6}\widehat{\rho}_{ij}\big)-2\sin\!\big(\tfrac{\pi}{6}\rho_{ij}\big)\right|\leq 8\pi\sqrt{\frac{\log p}{n}}.$
Regarding the term $|\widehat{B}_{ij}-B_{ij}|$, we define $f(a,b)=\dfrac{1+ab}{\sqrt{(1+a^{2})(1+b^{2})}}$ so that $\widehat{B}_{ij}=f(\widehat{\alpha}_{i},\widehat{\alpha}_{j})$ and $B_{ij}=f(\alpha_{i},\alpha_{j})$ . A direct calculation shows that:
$\frac{\partial f}{\partial a}(a,b)=\frac{b-a}{(1+a^{2})^{3/2}(1+b^{2})^{1/2}},\qquad\frac{\partial f}{\partial b}(a,b)=\frac{a-b}{(1+a^{2})^{1/2}(1+b^{2})^{3/2}},$
hence $|\partial f/\partial a|\leq 1$ and $|\partial f/\partial b|\leq 1$ on $\mathbb{R}^{2}$. By the mean value theorem in two variables,
$|\widehat{B}_{ij}-B_{ij}|=|f(\widehat{\alpha}_{i},\widehat{\alpha}_{j})-f(\alpha_{i},\alpha_{j})|\leq|\widehat{\alpha}_{i}-\alpha_{i}|+|\widehat{\alpha}_{j}-\alpha_{j}|.$ (20)
Therefore, we have that:
$\max_{ij}\left|\widehat{S}^{\rho,\alpha}_{ij}-{S}^{\rho,\alpha}_{ij}\right|\leq 8\pi\sqrt{\frac{\log p}{n}}+2\max_{1\leq l\leq p}|\widehat{\alpha}_{l}-\alpha_{l}|$
Define $A:=\max_{1\leq\ell\leq p}|\widehat{\alpha}_{\ell}-\alpha_{\ell}|$ and fix an arbitrary threshold $t_{\alpha}>0$. On the event $\{A\leq t_{\alpha}\}$ we have $2A\leq 2t_{\alpha}$. Hence:
$\mathbb{P}\big(2A\leq 2t_{\alpha}\big)\;\geq\;1-\mathbb{P}(A>t_{\alpha}).$
Equivalently, with probability at least $1-\mathbb{P}(A>t_{\alpha})$, the plug-in error term satisfies
$2\,\max_{1\leq\ell\leq p}\big|\widehat{\alpha}_{\ell}-\alpha_{\ell}\big|\;\leq\;2t_{\alpha}.$
Exploiting Boole’s inequality, we obtain:
$\mathbb{P}\!\left(\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;8\,\pi\,\sqrt{\frac{\log p}{n}}\;+\;2t_{\alpha}\right)\;\geq\;1-\frac{1}{p}-\mathbb{P}(A>t_{\alpha}).$
In particular, if $t_{\alpha}$ is chosen so that $\mathbb{P}(A>t_{\alpha})\leq 1/p^{2}$, then with probability at least $1-2/p^{2}$,
$\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;8\,\pi\,\sqrt{\frac{\log p}{n}}\;+\;2t_{\alpha}.$
Moreover, if we assume that the $\alpha$ parameters are correctly estimated using likelihood-based methods, we have that $A=O_{P}\!\left(\sqrt{\frac{\log p}{n}}\right)$, taking $t_{\alpha}=C_{\alpha}\sqrt{\frac{\log p}{n}}$, where $C_{\alpha}$ is a constant value that depends from $\alpha$, we obtain:
$\|\widehat{S}^{\rho,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;\big(8\,\pi+2C_{\alpha}\big)\sqrt{\frac{\log p}{n}}\quad,$
with probability greater than $1-\tfrac{2}{p^{2}}$.
@@
From liu2012high, we know that with probability $1-1/p^{2}$ we have that:
$\left|\widehat{S}_{ij}^{\tau}-\Sigma_{ij}^{0}\right|\leq 2.45\pi\sqrt{\frac{\log p}{n}}$
The results for our estimator become:
$\max_{ij}\left|\widehat{S}^{\tau,\alpha}_{ij}-{S}^{\tau,\alpha}_{ij}\right|\leq 2.45\pi\sqrt{\frac{\log p}{n}}+2\max_{1\leq l\leq p}|\widehat{\alpha}_{l}-\alpha_{l}|,$
where choosing a probability $1-1/p^{2}$, and assuming the convergence of the likelihood-based estimator, we obtain:
$\|\widehat{S}^{\tau,\alpha}-\Sigma^{0}\|_{\max}\;\leq\;\big(2.45\,\pi+2C_{\alpha}\big)\sqrt{\frac{\log p}{n}},$
with probability greater than $1-\tfrac{2}{p^{2}}$.